How does one prove that


(a/sqrt(a2+8bc)) + (b/sqrt(b2+8ac)) + (c/sqrt(c2+8ab)) >= 1 ?


sqrt() is the square root function

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mardava RJgpl

((a/sqrt(a2+8bc)) + (b/sqrt(b2+8ac)) + (c/sqrt(c2+8ab)) >= 1

 

USe AM >=GM

((a/sqrt(a2+8bc)) + (b/sqrt(b2+8ac)) + (c/sqrt(c2+8ab))/3 >=((a/sqrt(a2+8bc)) (b/sqrt(b2+8ac))  (c/sqrt(c2+8ab))1/3    ............................1

 

  Now use again AM>= GM

Let three number  a2,4bc,4bc

  (a2 +4bc +4bc)/3 >=(16a2b2c2)1/3

  (a2 +8bc ) >=3(16a2b2c2)1/3

or sqrt(a2 +8bc )>=√3(4abc)1/3

use this result in equation 1

((a/sqrt(a2+8bc)) + (b/sqrt(b2+8ac)) + (c/sqrt(c2+8ab)) >=3{abc/(33/24abc)}1/3

 

or  

((a/sqrt(a2+8bc)) + (b/sqrt(b2+8ac)) + (c/sqrt(c2+8ab)) >=√3/41/3

                                                                                                            >=1.09

                                                                                                             >1


Please feel free to post as many doubts on our discussion forum as you can. If you find any

question Difficult to understand - post it here and we will get you the answer and detailed

solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin